I want to know how to factor from the standard form of a quadratic equation. I'm in the beginning of college algebra, and can't remember how to do it.
The problem is: 10x^2 - 3x - 4 = 0
how do i figure out x without doing the whole quadratic equation?
I know the answer to factoring is (2x + 1) (5x - 4). how do you get that??

Answer by vasu2qute(17)

Given 10x^2 - 3x - 4 = 0
10x^2 -8x + 5x - 4 = 0
2x(5x - 4) + 1(5x - 4) = 0
(5x-4)(2x+1) = 0
(I have written -3x as (-8x+5x) then took 2x common from the first two terms and 1 common for the next two terms . In the next step I took 5x-4 as common)